# LP-ICP: General Localizability-Aware Point Cloud Registration for Robust Localization in Extreme Unstructured Environments

### v2

@@ VI-A2 Algorithmic Comparisons
We integrate the current state-of-the-art methods Zhang et al. [20] and X-ICP [21] into the point cloud registration framework of LVI-SAM for experiments and comparisons. In LVI-SAM, the degeneracy detection method used is based on Zhang et al. [20]. To ensure that [20] performs well for degeneracy detection and handling in all experiments, we empirically set the threshold of Zhang et al. [20] to 50. We also conducted experiments to evaluate its performance with different threshold settings, as shown in Table I. Finally, we choose a threshold of 50 for the subsequent experimental evaluations. Additionally, since the source code for X-ICP [21] is not yet publicly available, we reproduced the method based on the paper and integrated it into the point cloud registration framework of LVI-SAM. However, during our experiments, adding the hard constraints in the partially locatable directions, as described in [21], caused large fluctuations in pose updates, often resulting in localization failures. Therefore, we modified these partially locatable directions to be non-locatable for the experiment. The other threshold values used are based on the original threshold settings in [21]. In the experiments, we refer to it as Xn-ICP. Moreover, when implementing Xn-ICP in LVI-SAM, since X-ICP [21] is designed for point-to-plane ICP, we only used point-to-plane correspondences for localizability detection. However, for fairness, both edge point-to-line and planar point-to-plane correspondences were used during optimization to retain as much usable information as possible for the optimization.
@@
In addition, we also tested the performance of other advanced SLAM systems in our experiments for comparison with the accuracy of our method.
@@ VI-B Simulation Experiments in a Planetary-like Environment
Figure 6: Trajectories estimated by LVI-SAM + Ours and the state-of-the-art methods on the sequences a1_home, a3_odom, a2_traverse, a4_home, a7_loop, and a3_final, and ground truth trajectories are shown.
@@
VI-B1 PLAM Dataset
VI-B2 Results
We evaluate the localization performance of LVI-SAM + Ours, LVI-SAM + Zhang et al. [20], and LVI-SAM + Xn-ICP [21] on the PLAM dataset and compare and analyze the results. At the same time, we test and evaluate the performance of some advanced SLAM systems, such as ORB-SLAM3 [43] in monocular camera-IMU mode, VINS-MONO [44], and LIO-SAM [31]. The estimated trajectory obtained by each algorithm with the ground truth trajectory are compared. The first 50 poses of both trajectories are aligned using the EVO tool [46], and the Absolute Trajectory Error (ATE) [47] between them is calculated. The results are presented in the form of root mean square error (RMSE). Additionally, we provide an extra metric, which is the ratio of the algorithm’s estimated poses to the total length of the ground truth for each sequence. This value is calculated as the length of the ground truth associated with the estimated trajectory divided by the total length of the ground truth for the sequence. When large drift or fluctuations in the trajectory occurred, preventing the algorithm from continuing, we stopped the experiment. Table II reports these metrics for the tested algorithms on each sequence of the dataset. These results show that the extreme environment of the PLAM dataset poses significant challenges to VINS-MONO, ORB-SLAM3, and LIO-SAM. The long-term visual ambiguity, lack of features and geometric structure, and poor lighting conditions in the PLAM dataset lead to instability and low accuracy during their operation. In comparison, the multi-sensor fusion algorithm LVI-SAM achieves better accuracy than them.
The degeneracy detection algorithms in LVI-SAM + Ours, LVI-SAM + Zhang et al. [20], and LVI-SAM + Xn-ICP [21] were further compared. The trajectories estimated by these three methods on six sequences of the dataset are presented in Fig. 6. Our method, LP-ICP, utilizes localizability information from edge point-to-line and planar point-to-plane correspondences to detect degeneracy, achieving the best localization accuracy across all sequences. LVI-SAM + Zhang et al. [20] also demonstrates good accuracy in the rocky-area sequences a2_home, a4_loop, and a7_loop, as the abundance of rocks provides more features and geometric structures. However, in sequences from other non-rocky areas, on the one hand, the method cannot pick out high-contribution correspondence constraints in ill-conditioned directions, thus failing to utilize more effective information in binary degeneracy detection. On the other hand, the scales of rotational and translational directions differ. If the same threshold settings are applied, it limits the degeneracy detection, leading to a decrease in performance. For instance, as shown in Fig. 7, in the a2_traverse sequence, when the robot is in an open area of the scene at 620s, LiDAR degeneracy occurs in the three directions: translation along the x-axis and y-axis, and rotation around the z-axis. However, the method by Zhang et al. [20] can only detect two of these directions. X-ICP improves performance by adding partial localizability categories, which allows it to utilize more constraint information. It also unifies the scale of localizability for both rotation and translation. In the a2_odom and a4_home sequences, it shows better accuracy compared to Zhang et al. [20]. However, X-ICP cannot utilize the localizability information from edge point-to-line correspondences. As shown in Fig. 7, compared to our method, it has more instances of non-localizable categories. This prevents further improvement in its localization accuracy. Additionally, in partially localizable directions, it adds hard constraints to the optimization process. This is sensitive to the quality of resampled points and increases the risk of significant trajectory drift and fluctuations, which can lead to localization failure. This is particularly problematic for long-range sequences or sequences without loops in non-rock areas, such as a2_traverse and a3_odom. Although, for the purpose of comparison, when reproducing X-ICP, we handled the optimization in partially localizable directions as non-localizable directions. This avoided direct localization failure. However, its accuracy remains unsatisfactory on some challenging sequences.
@@ VI-C Real-World Experiment on the CERBERUS DARPA Subterranean Challenge Dataset
As shown in Table III, the characteristics of underground tunnel pose significant challenges to many methods. Poor lighting conditions and darkness, along with self-similar scenes, create difficulties for vision-based approaches. LiDAR -based methods struggle with accuracy in degenerate scenarios, such as long corridor-like passages in tunnels. Uneven terrain and vibrations during movement also negatively impact the final localization results. These factors result in unsatisfactory accuracy for VINS-MONO, LIO-SAM, and R3LIVE. By employing multi-sensor fusion and incorporating degeneracy detection and handling, LVI-SAM + Ours, LVI-SAM + Zhang et al., and LVI-SAM + Xn-ICP achieve significantly better localization accuracy. Among these three localizability-aware algorithms, our method achieves the best accuracy on the ANYmal 1, ANYmal 2, and ANYmal 3 sequences. On the ANYmal 4 sequence, its accuracy is comparable to that of LVI-SAM + Zhang et al.
@@
Additionally, the mapping results of the three methods on the ANYmal 1 sequence are shown in Fig. 8. The magnified parts show the most accurate map to be produced by the proposed method. In the dashed box of Fig. 8(b), which marks the tunnel connection area A, the robot approached A from both ends during operation but ultimately did not reach this location. This posed a challenge for mapping. As shown in the mapping result (b) of Xn-ICP, which presents the top-down view in the x-y plane, there are noticeable disconnections in the map at location A. As shown in the mapping result (c) of Zhang et al. [20], which presents the front view in the x-z plane at the same location, the mapping result contains an error where the top of the tunnel on the right side of A is lower than the ground of the left side tunnel. This causes an incorrect tunnel connection at location A. In contrast, the result of our method, LP-ICP, does not exhibit these errors.
VI-D Runtime Evaluation
@@
Table IV shows the runtime results of LP-ICP on the ANYmal 1 sequence of the CERBERUS DARPA subterranean challenge dataset. We tested the average time consumption of the scan-to-map registration module for three methods: LVI-SAM (without degeneracy detection), LVI-SAM + Ours, and LVI-SAM + Zhang et al [20]. The tests were conducted on a laptop with an Intel i7-12700H CPU. As shown in Table IV, the time consumption of the three methods is quite similar. Interestingly, the method without degeneracy detection takes the most time, which we believe is due to optimization in ill-conditioned directions that makes it harder to converge, resulting in an increased number of iterations. This demonstrates that our method, LP-ICP, has the capability to run in real time on robotic systems.
@@
TABLE IV: The Average Time Consumption (Milliseconds) of Scan-to-Map Registration Using Different Methods on the Sequence ANYmal 1
@@ VII Conclusion
To improve the accuracy of LiDAR-based SLAM algorithms in challenging unstructured environments, degeneracy detection and handling are essential. This paper proposes a new ICP algorithm framework, LP-ICP, which combines point-to-line and point-to-plane cost functions, along with localizability detection and handling. LP-ICP detects degeneracy by calculating the localizability contribution of geometric correspondences and adds additional constraints to improve the accuracy of pose estimation. The efficacy of the proposed method is validated through experimental evaluation on our planetary-like simulation dataset and a real-world underground tunnel dataset, demonstrating higher or comparable accuracy than the state-of-the-art methods.

### v3

@@ VI-A2 Algorithmic Comparisons
We integrate the current state-of-the-art methods Zhang et al. [21] and X-ICP [22] into the point cloud registration framework of LVI-SAM for experiments and comparisons. In LVI-SAM, the degeneracy detection method used is based on Zhang et al. [21]. To ensure that [21] performs well for degeneracy detection and handling in all experiments, we empirically set the threshold of Zhang et al. [21] to 50. We also conducted experiments to evaluate its performance with different threshold settings, as shown in Table I. Finally, we choose a threshold of 50 for the subsequent experimental evaluations. Additionally, we integrated X-ICP [22] into the point cloud registration framework of LVI-SAM, based on the paper and its open-source implementation [^1] https://github.com/leggedrobotics/perfectlyconstrained. After parameter tuning, the thresholds were empirically set to $\kappa_{1}=90$, $\kappa_{2}=50$, and $\kappa_{3}=35$. In our experiments, the version integrated into LVI-SAM is referred to as Xn-ICP. In addition, we also evaluated the performance of several advanced SLAM systems in our experiments.
@@ VI-B1 PLAM Dataset
Figure 6: (a)-(f) Trajectories estimated by LVI-SAM + Ours and the state-of-the-art methods on the sequences a1_home, a3_odom, a2_traverse, a4_home, a7_loop, and a3_final, and ground truth trajectories are shown. (g) Edge points (green) and planar points (purple) extracted by the front end of LVI-SAM from a LiDAR scan during the a1_home sequence.
VI-B2 Results
We evaluate the localization performance of LVI-SAM + Ours, LVI-SAM + Zhang et al. [21], and LVI-SAM + Xn-ICP [22] on the PLAM dataset and compare and analyze the results. At the same time, we test and evaluate the performance of some advanced SLAM systems, such as ORB-SLAM3 [49] in monocular camera-IMU mode, VINS-MONO [48], and LIO-SAM [33]. The estimated trajectory obtained by each algorithm with the ground truth trajectory are compared. The first 50 poses of both trajectories are aligned using the EVO tool [50], and the Absolute Trajectory Error (ATE) [51] between them is calculated. The results are presented in the form of root mean square error (RMSE). Additionally, we provide an extra metric, which is the ratio of the algorithm’s estimated poses to the total length of the ground truth for each sequence. This value is calculated as the length of the ground truth associated with the estimated trajectory divided by the total length of the ground truth for the sequence. When large drift or fluctuations in the trajectory occurred, preventing the algorithm from continuing, we stopped the experiment. We compare the accuracy among algorithms that successfully complete the entire sequence. For those that did not complete the sequence, we provided their ATE to better understand their performance on the specific sequence. Table II reports these metrics for the tested algorithms on each sequence of the dataset. These results show that the extreme environment of the PLAM dataset poses significant challenges to VINS-MONO, ORB-SLAM3, and LIO-SAM. The long-term visual ambiguity, lack of features and geometric structure, and poor lighting conditions in the PLAM dataset lead to instability and low accuracy during their operation. In comparison, the multi-sensor fusion algorithm LVI-SAM achieves better accuracy than them.
The degeneracy detection algorithms in LVI-SAM + Ours, LVI-SAM + Zhang et al. [21], and LVI-SAM + Xn-ICP [22] were further compared. The trajectories estimated by these three methods on six sequences of the dataset are presented in Fig. 6. Fig. 6 (g) shows the edge points and planar points extracted from a LiDAR scan during the a1_home sequence, when degeneracy is detected along a translational direction in the x-y plane. Our method, LP-ICP, achieves better localization accuracy on multiple sequences. LVI-SAM + Zhang et al. [21] also demonstrates good accuracy in the rocky-area sequences a2_home, a4_loop, and a7_loop, as the abundance of rocks provides more features and geometric structures. However, in sequences from other non-rocky areas, on the one hand, the method cannot pick out high-contribution correspondence constraints in ill-conditioned directions. On the other hand, the scales of rotational and translational directions differ. If the same threshold settings are applied, it limits the degeneracy detection, leading to a decrease in performance. For instance, as shown in Fig. 7, in the a2_traverse sequence, when the robot is in an open area of the scene at 620s, LiDAR degeneracy occurs in the three directions: translation along the x-axis and y-axis, and rotation around the z-axis. However, the method by Zhang et al. [21] can only detect two of these directions. X-ICP improves performance by adding partial localizability categories and unifying the scale of localizability for both rotation and translation. In multiple sequences such as a2_odom and a4_home, it shows better accuracy compared to Zhang et al. [21]. However, X-ICP cannot utilize the localizability information from edge point-to-line correspondences. As shown in Fig. 7, our method considers more localizability information to be usable. Additionally, in partially localizable directions, the added soft constraints enable pose updates to be performed under additional constraints and help improve accuracy. However, in our experiments, we observed that the use of soft constraints in partially localizable directions may introduce fluctuations and uncertainty in pose estimation. The underlying causes and potential improvements will be investigated in future work.
@@ VI-C Real-World Experiment on the CERBERUS DARPA Subterranean Challenge Dataset
As shown in Table III, the characteristics of underground tunnel pose significant challenges to many methods. Poor lighting conditions and darkness, along with self-similar scenes, create difficulties for vision-based approaches. LiDAR -based methods struggle with accuracy in degenerate scenarios, such as long corridor-like passages in tunnels. Uneven terrain and vibrations during movement also negatively impact the final localization results. These factors result in unsatisfactory accuracy for VINS-MONO, LIO-SAM, and R3LIVE. By employing multi-sensor fusion and utilizing the VIO submodule to provide a relatively reliable initial estimate, along with incorporating degeneracy detection and handling, LVI-SAM + Ours, LVI-SAM + Zhang et al., and LVI-SAM + Xn-ICP achieve significantly better localization accuracy. Among these three localizability-aware algorithms, our method achieves better accuracy on the ANYmal 2 and ANYmal 3 sequences. On the ANYmal 1 and ANYmal 4 sequences, its accuracy is comparable to that of LVI-SAM + Xn-ICP and LVI-SAM + Zhang et al.
@@
Additionally, the mapping results of the three methods on the ANYmal 1 sequence are shown in Fig. 8. In the dashed box of Fig. 8, which marks the tunnel connection area A, the robot approached A from both ends during operation but ultimately did not reach this location. This posed a challenge for mapping. As shown in the mapping result of Zhang et al. [21], which presents the front view of location A in the x-z plane, the mapping result contains an error where the top of the tunnel on the right side of A is lower than the ground of the left side tunnel. This causes an incorrect tunnel connection at location A. In contrast, by using fine-grained degeneracy detection methods, the results of LVI-SAM + Ours and LVI-SAM + Xn-ICP do not exhibit this error.
Figure 9: (a) Trajectories estimated by three methods and ground truth trajectory are shown. (b) Estimated localizability categories of LP-ICP and the state-of-the-art methods. (c) Edge points (green) and planar points (purple) extracted from three LiDAR scans collected at locations A, B, and C in (a), respectively.
VI-D SubT-MRS Dataset
The SubT-MRS Dataset [54] is a challenging real-world dataset collected from subterranean environments, including caves, urban areas, and tunnels. It was used in the ICCV 2023 SLAM Challenge. The dataset introduces challenging conditions such as textureless surfaces, smoke, illumination changes, and featureless structures.
We conducted tests on the Long Corridor sequence from the SubT-MRS dataset. In this scenario, the localization accuracy of the ground robot is primarily affected by degeneracy along the longitudinal translational direction of the corridor. In this experiment, camera images were not used, and loop closure detection was disabled. Since LVI-SAM becomes similar to LIO-SAM when image topics are not subscribed to, the three compared methods are referred to as LIO-SAM + Zhang et al., LIO-SAM + Xn-ICP, and LIO-SAM + Ours.
Table IV presents the localization accuracy of the three compared methods. The estimated trajectories of the three methods are shown in Fig. 9 (a). The localization errors are primarily caused by drift at two specific locations. Fig. 9 (b) shows the degeneracy detection results of the three methods along the translational direction with the weakest constraints. As shown in Fig. 9(a), all three methods begin to exhibit noticeable drift after reaching position A from the start point. This is due to the lack of sufficient information from edge or planar features along the corridor direction, as shown in Fig. 9 (c) A. Our method incorporates additional point-to-line localizability information near location B, resulting in reduced drift. Note that between 130s and 170s, when the robot approaches the end of the corridor, LIO-SAM + Xn-ICP detects degeneracy along the translational direction. We attribute this to drift-induced mapping errors, which result in incorrect scan-to-map matching and erroneous normal vector estimation in the front end. This issue is not inherent to X-ICP itself. LIO-SAM + Xn-ICP exhibits drift near position C. In contrast, LIO-SAM + Zhang et al. and LIO-SAM + Ours utilize edge point-to-line geometric constraints at this location, as shown in Fig. 9 (c) C.
TABLE IV: The RMSE of ATE of the Compared Algorithms on the Long Corridor sequence from the SubT-MRS dataset
ATE [m]
LIO-SAM + Zhang et al. [21] 24.71
LIO-SAM + Xn-ICP [22] 19.63
LIO-SAM + Ours 11.92
  • Bold value indicates the best result.
VI-E Runtime Evaluation
@@
Table V shows the runtime results of LP-ICP on the ANYmal 1 sequence of the CERBERUS DARPA subterranean challenge dataset. We tested the average time consumption of the scan-to-map registration module for three methods: LVI-SAM (without degeneracy detection), LVI-SAM + Ours, and LVI-SAM + Zhang et al [21]. The tests were conducted on a laptop with an Intel i7-12700H CPU. As shown in Table V, the time consumption of the three methods is quite similar. Interestingly, the method without degeneracy detection takes the most time, which we believe is due to optimization in ill-conditioned directions that makes it harder to converge, resulting in an increased number of iterations. This demonstrates that our method, LP-ICP, has the capability to run in real time on robotic systems.
@@
TABLE V: The Average Time Consumption (Milliseconds) of Scan-to-Map Registration Using Different Methods on the ANYmal 1 Sequence
@@ VII Conclusion
To improve the accuracy of LiDAR-based SLAM algorithms in challenging unstructured environments, degeneracy detection and handling are essential. This paper proposes a new ICP algorithm framework, LP-ICP, which combines point-to-line and point-to-plane cost functions, along with localizability detection and handling. LP-ICP detects degeneracy by calculating the localizability contribution of geometric correspondences and adds additional constraints to improve the accuracy of pose estimation. The efficacy of the proposed method is validated through experimental evaluation on our planetary-like simulation dataset and two real-world datasets, demonstrating comparable or improved performance over state-of-the-art methods under our tested settings. Currently, the performance of degeneracy detection and optimization algorithms is influenced by their parameter settings. In our future work, we will validate the sensitivity of the algorithm parameters to different environments and hardware configurations, as well as investigate the impact of performing localizability analysis separately in the translational and rotational eigenspaces. In addition, the partial localizable category with soft constraints introduces trajectory jitter and uncertainty in the experimental results. In future work, we will further investigate the underlying causes and explore potential improvements.